# Splittings of One-Ended Groups with One-Ended Halfspaces

### v3

@@ 4.4 The pocset $\mathcal{P}$
      Finally, if $(\mathfrak{a}_{2},\mathfrak{h}_{2})=(g[z]^{*},g\mathfrak{h}_{0}^{*})$, with $g\in G$ and $[z]\in\mathcal{M}_{0}$, then $[x]\subseteq\mathfrak{h}_{0}\subseteq\mathfrak{h}_{2}\preceq g[z]^{*}$, so the first row of (9) holds (and they are both inequalities of type (2)).∎
@@ 4.5 The tree $T^{\prime}$
Let $T^{\prime}$ be the cubing of $\mathcal{P}$. Lemma 4.20 tells us that $T^{\prime}$ is a tree. Moreover, the action of $G$ on $\mathcal{P}$ induces an action of $G$ on $T^{\prime}$. The edges in $T^{\prime}$ correspond to pairs $\{(\mathfrak{a},\mathfrak{h}),(\mathfrak{a},\mathfrak{h})^{*}\}\subseteq\mathcal{P}$, and, since $(\mathfrak{a},\mathfrak{h})^{*}$ always takes the form $(\mathfrak{a}^{\prime},\mathfrak{h}^{*})$, an edge inversion in $T^{\prime}$ would mean there is an element $g\in G$ with $g\mathfrak{h}=\mathfrak{h}^{*}$. But that would give rise to an edge inversion in $T$, which doesn’t happen, hence $G$ acts on $T^{\prime}$ without edge inversions, i.e. $G\curvearrowright T^{\prime}$ is a splitting. We may assume that the action $G\curvearrowright T^{\prime}$ is minimal since we can always replace $T^{\prime}$ with a minimal $G$-invariant subtree. The description of edges in $T^{\prime}$ also makes it clear that every edge stabilizer of $T^{\prime}$ is contained in an edge stabilizer of $T$.
@@
$\phi(v):=$$\{\mathfrak{h}\in\mathcal{H}^{\#}\mid(\mathfrak{h},\mathfrak{h})\in v\}$
@@
Let’s first prove completeness of $\phi(v)$. For $\mathfrak{h}\in\mathcal{H}^{\#}$, exactly one of $(\mathfrak{h},\mathfrak{h})$ and $(\mathfrak{h}^{*},\mathfrak{h}^{*})$ is in $v$ by completeness of $v$, so exactly one of $\mathfrak{h}$ and $\mathfrak{h}^{*}$ is in $\phi(v)$. For $g\in G$ and $[x]\in\mathcal{M}_{0}$, exactly one of $(g[x],g\mathfrak{h}_{0})$ and $(g[x]^{*},g\mathfrak{h}_{0}^{*})$ is in $v$ by completeness of $v$, so for each $g\in G$ we have exactly one of $g\mathfrak{h}_{0}$ and $g\mathfrak{h}_{0}^{*}$ in $\phi(v)$.
Before showing that $\phi(v)$ is consistent and DCC, we make the following observation, which follows from Lemmas 4.13 and 4.14: if $\mathfrak{h}_{1}\subsetneq\mathfrak{h}_{2}$ in $\mathcal{H}(X)$ then
@@
    if $\mathfrak{h}_{1}\in\mathcal{H}^{\#}$ then $\mathfrak{a}_{1}=\mathfrak{h}_{1}$,
@@
    if $\mathfrak{h}_{1}=g_{1}\mathfrak{h}_{0}$ then $\mathfrak{a}_{1}=g_{1}[x]$ and (13) holds for all choices of $[x]\in\mathcal{M}_{0}$,
@@
    if $\mathfrak{h}_{1}=g_{1}\mathfrak{h}_{0}^{*}$ then $\mathfrak{a}_{1}=g_{1}[x]^{*}$ for the unique $[x]\in\mathcal{M}_{0}$ with $\mathfrak{h}_{2}^{*}\subseteq g_{1}[x]$,
@@
    if $\mathfrak{h}_{2}\in\mathcal{H}^{\#}$ then $\mathfrak{a}_{2}=\mathfrak{h}_{2}$,
@@
    if $\mathfrak{h}_{2}=g_{2}\mathfrak{h}_{0}$ then $\mathfrak{a}_{2}=g_{2}[y]$ for the unique $[y]\in\mathcal{M}_{0}$ with $\mathfrak{h}_{1}\subseteq g_{2}[y]$,
@@
    if $\mathfrak{h}_{2}=g_{2}\mathfrak{h}_{0}^{*}$ then $\mathfrak{a}_{2}=g_{2}[y]^{*}$ and (13) holds for all choices of $[y]\in\mathcal{M}_{0}$.
We now prove consistency of $\phi(v)$. If $\mathfrak{h}_{1}\subsetneq\mathfrak{h}_{2}$ and $\mathfrak{h}_{1}\in\phi(v)$ then $(\mathfrak{a}_{1},\mathfrak{h}_{1})\in v$, where $(\mathfrak{a}_{1},\mathfrak{h}_{1})$ is given by the appropriate case from (1)(3) (and in case (2) we only have $(g_{1}[x],g_{1}\mathfrak{h}_{0})\in v$ for one $[x]\in\mathcal{M}_{0}$). Consistency of $v$ then implies that $(\mathfrak{a}_{2},\mathfrak{h}_{2})\in v$, where $(\mathfrak{a}_{2},\mathfrak{h}_{2})$ is given by the appropriate case from (4)(6) (and in case (6) we have $(g_{2}[y]^{*},g_{2}\mathfrak{h}_{0}^{*})\in v$ for all $[y]\in\mathcal{M}_{0}$). It follows that $\mathfrak{h}_{2}\in\phi(v)$.
Finally, we prove that $\phi(v)$ is DCC. Suppose for contradiction that we have an infinite descending chain $\mathfrak{h}_{1}\supsetneq\mathfrak{h}_{2}\supsetneq\mathfrak{h}_{3}\supsetneq\dots$ in $\phi(v)$. Then we can turn this into an infinite descending chain
@@
    if $\mathfrak{h}_{i}\in\mathcal{H}^{\#}$ then $\mathfrak{a}_{i}=\mathfrak{h}_{i}$,
@@
    if $\mathfrak{h}_{i}=g\mathfrak{h}_{0}$ then $\mathfrak{a}_{i}=g[x]$ for the unique $[x]\in\mathcal{M}_{0}$ with $\mathfrak{h}_{i+1}\subseteq g[x]$,
@@
    if $\mathfrak{h}_{i}=g\mathfrak{h}_{0}^{*}$ then $\mathfrak{a}_{i}=g[x]^{*}$ for the unique $[x]\in\mathcal{M}_{0}$ with $\mathfrak{h}_{i-1}^{*}\subseteq g[x]$ (or any $[x]\in\mathcal{M}_{0}$ if $i=1$).
@@
Recall that we have a wallspace $(V\mathfrak{h}_{0},\mathcal{P}_{0})$ with cubing $T_{0}$, and an associated map $\lambda:V\mathfrak{h}_{0}\to VT_{0}$. Moreover, it follows from the constructions that $\lambda$ is $G_{0}$-equivariant. We observed that the $\sim$-classes $[x]$ are precisely the non-empty fibers of $\lambda$, so in particular we get a well-defined injective $G_{0}$-equivariant map $\mathcal{M}_{0}\to VT_{0}$ given by $[x]\mapsto\lambda(x)$.
@@
  • For $\mathfrak{h}\in\mathcal{H}^{\#}$, we have an edge $e^{\prime}=\{(\mathfrak{h},\mathfrak{h}),(\mathfrak{h}^{*},\mathfrak{h}^{*})\}\in ET^{\prime}$ and also an edge $e\in ET$. In this case $e\notin G\cdot e_{0}$, so the tree $T_{e}$ is just a single vertex, and we define $\tau(e^{\prime})$ to be this vertex. The maps #$\mathcal{H}^{\#}\to ET$ and #$\mathcal{H}^{\#}\to ET^{\prime}$ are $G$-equivariant, so $\tau(ge^{\prime})=g\cdot\tau(e^{\prime})$ for all $g\in G$.
@@
    with $i\in\Omega$ and $[x]\in\mathcal{M}_{0}$. We then define
@@
Finally, the map $\tau:ET^{\prime}\to\sqcup_{e}VT_{e}$ is injective because the map $\mathcal{M}_{0}\to VT_{0}$ given by $[x]\mapsto\lambda(x)$ is injective. ∎
@@
Recall that $e_{0}\in ET$ is the edge corresponding to the halfspace $\mathfrak{h}_{0}$, in particular $G_{0}=G_{e_{0}}$. The vertex $v_{0}\in VT$ will be the endpoint of $e_{0}$ that is contained in the halfspace $p(\mathfrak{h}_{0})\in\mathcal{H}(T)$. The vertex stabilizer $G_{v_{0}}$ fixes no edges in $T$, so there exists $g\in G_{v_{0}}-G_{e_{0}}$. Then $g\mathfrak{h}_{0}^{*}\subsetneq\mathfrak{h}_{0}$. By Lemma 4.14, there exist $[x],[y]\in\mathcal{M}_{0}$ with $g[y]^{*}\preceq[x]$, so $(g[y]^{*},g\mathfrak{h}_{0}^{*})\leq([x],\mathfrak{h}_{0})$ in $\mathcal{P}$. Recall that we have a $G_{0}$-equivariant map $\mathcal{M}_{0}\to T_{0}$ (as in the proof of Lemma 4.22), and $G_{0}$ has no fixed point in $T_{0}$ by Lemma 4.12, so the $G_{0}$-orbits in $\mathcal{M}_{0}$ are all infinite. In particular, there exists $g_{0}\in G_{0}$ with $g_{0}[x]\neq[y]$. Then $(g_{0}[x],\mathfrak{h}_{0})\leq([y]^{*},\mathfrak{h}_{0}^{*})$ by Lemma 4.17. Putting this all together, we get

### v4

@@ 4.6 The pocset $\mathcal{P}$
      Finally, if $(\mathfrak{a}_{2},\mathfrak{h}_{2})=(g[z]^{*},g\mathfrak{h}_{0}^{*})$, with $g\in G$ and $z]\in\mathcal{M}_{0}$, then $x]\subseteq\mathfrak{h}_{0}\subseteq\mathfrak{h}_{2}\preceq g[z]^{*}$, so the first row of (12) holds (and they are both inequalities of type (2)).∎
In the next four lemmas we prove that $\mathcal{P}$ is a discrete pocset (see Definition 3).
Lemma 4.22.
If $(\mathfrak{a}_{1},\mathfrak{h}_{1})\leq(\mathfrak{a}_{2},\mathfrak{h}_{2})$ with $\hat{\mathfrak{h}}_{1}\neq\hat{\mathfrak{h}}_{2}$ then $V\hat{\mathfrak{h}}_{1}\subseteq\mathfrak{a}_{2}$.
Proof.
We split into three cases depending on the form of $(\mathfrak{a}_{2},\mathfrak{h}_{2})$.
  1. (1)
    Suppose $(\mathfrak{a}_{2},\mathfrak{h}_{2})=(V\mathfrak{h}_{2},\mathfrak{h}_{2})$ with $\mathfrak{h}_{2}\in\mathcal{H}^{\#}$. If $\hat{\mathfrak{h}}_{1}\subseteq\mathfrak{h}_{2}$ then we are done, so suppose $\hat{\mathfrak{h}}_{1}\subseteq\mathfrak{h}_{2}^{*}$. As $\mathfrak{a}_{1}\preceq\mathfrak{a}_{2}=V\mathfrak{h}_{2}$ and $\hat{\mathfrak{h}}_{1}\neq\hat{\mathfrak{h}}_{2}$, we reduce to two possibilities for $(\mathfrak{a}_{1},\mathfrak{h}_{1})$:
    1. (a)
      Suppose $(\mathfrak{a}_{1},\mathfrak{h}_{1})=(V\mathfrak{h}_{1},\mathfrak{h}_{1})$ with $\mathfrak{h}_{2}\subsetneq\mathfrak{h}_{1}$. Then $\mathfrak{a}_{2}=V\mathfrak{h}_{2}\subseteq V\mathfrak{h}_{1}=\mathfrak{a}_{1}$ and $\mathfrak{h}_{1}\not\subseteq\mathfrak{h}_{2}$, contradicting $(\mathfrak{a}_{1},\mathfrak{h}_{1})\leq(\mathfrak{a}_{2},\mathfrak{h}_{2})$.
    2. (b)
      Suppose $(\mathfrak{a}_{1},\mathfrak{h}_{1})=(g[x],g\mathfrak{h}_{0})$ with $\mathfrak{h}_{2}\subsetneq g\mathfrak{h}_{0}$. By Lemma 4.14 we know that $V\mathfrak{h}_{2}\subseteq g[y]$ for some $y]\in\mathcal{M}_{0}$. If $x]\neq[y]$ then $g[x]=\mathfrak{a}_{1}\preceq\mathfrak{a}_{2}=V\mathfrak{h}_{2}\subseteq g[y]\subseteq g[x]^{*}$, contradicting Lemma 4.16. If $[x]=[y]$ then $\mathfrak{a}_{2}=V\mathfrak{h}_{2}\subseteq g[y]=\mathfrak{a}_{1}$ and we contradict $(\mathfrak{a}_{1},\mathfrak{h}_{1})\leq(\mathfrak{a}_{2},\mathfrak{h}_{2})$ (noting $\mathfrak{h}_{1}\not\subseteq\mathfrak{h}_{2}$).
  2. (2)
    Suppose $(\mathfrak{a}_{2},\mathfrak{h}_{2})=(g_{2}[x_{2}],g_{2}\mathfrak{h}_{0})$. If $\hat{\mathfrak{h}}_{1}\subseteq\mathfrak{h}_{2}^{*}$ then we can obtain a contradiction by the same argument as in case (1), so we may suppose $\hat{\mathfrak{h}}_{1}\subseteq\mathfrak{h}_{2}=g_{2}\mathfrak{h}_{0}$. By Lemma 4.14 we have $V\hat{\mathfrak{h}}_{1}\subseteq g_{2}[y_{2}]$ for some $y_{2}]\in\mathcal{M}_{0}$. If $x_{2}]=[y_{2}]$ we are done, so suppose $x_{2}]\neq[y_{2}]$.
    1. (a)
      Suppose $\mathfrak{h}_{1}\subseteq g_{2}\mathfrak{h}_{0}$. Then $V\mathfrak{h}_{1}\subseteq g_{2}[y_{2}]$, which is disjoint from $\mathfrak{a}_{2}=g_{2}[x_{2}]$, so any path from $\mathfrak{a}_{1}\cap\mathfrak{h}_{1}$ to $\mathfrak{a}_{2}=g_{2}[x_{2}]$ must first cross $\hat{\mathfrak{h}}_{1}$. But $\mathfrak{a}_{1}\cap\mathfrak{h}_{1}\not\preceq\hat{\mathfrak{h}}_{1}$ (indeed this holds for any $(\mathfrak{a},\mathfrak{h})\in\mathcal{P}$), so $\mathfrak{a}_{1}\not\preceq\mathfrak{a}_{2}$, contradicting $(\mathfrak{a}_{1},\mathfrak{h}_{1})\leq(\mathfrak{a}_{2},\mathfrak{h}_{2})$.
    2. (b)
      Suppose $\mathfrak{h}_{1}^{*}\subseteq g_{2}\mathfrak{h}_{0}$. If $V\mathfrak{h}_{2}^{*}=g_{2}V\mathfrak{h}_{0}^{*}\subseteq\mathfrak{a}_{1}$ then we contradict $\mathfrak{a}_{1}\preceq\mathfrak{a}_{2}$, hence $(\mathfrak{a}_{1},\mathfrak{h}_{1})$ must take the form $(g_{1}[x_{1}],g_{1}\mathfrak{h}_{0})$, with $g_{2}V\mathfrak{h}_{0}^{*}\not\subseteq g_{1}[x_{1}]$. By Lemma 4.14 we have $g_{2}V\mathfrak{h}_{0}^{*}\subseteq g_{1}[y_{1}]$ for some $x_{1}]\neq[y_{1}]\in\mathcal{M}_{0}$. But then Lemma 4.15 implies that $g_{2}[y_{2}]^{*}\preceq g_{1}[y_{1}]$, so $g_{1}[x_{1}]=\mathfrak{a}_{1}\preceq\mathfrak{a}_{2}=g_{2}[x_{2}]\subseteq g_{2}[y_{2}]^{*}\preceq g_{1}[y_{1}]\subseteq g_{1}[x_{1}]^{*}$. This contradicts Lemma 4.16.
  3. (3)
    Suppose $(\mathfrak{a}_{2},\mathfrak{h}_{2})=(g_{2}[x_{2}]^{*},g_{2}\mathfrak{h}_{0}^{*})$. If $\hat{\mathfrak{h}}_{1}\subseteq\mathfrak{h}_{2}=g_{2}\mathfrak{h}_{0}^{*}$ then we are done, so suppose $\hat{\mathfrak{h}}_{1}\subseteq g_{2}\mathfrak{h}_{0}$. By Lemma 4.14 we have $V\hat{\mathfrak{h}}_{1}\subseteq g_{2}[y_{2}]$ for some $y_{2}]\in\mathcal{M}_{0}$. If $x_{2}]\neq[y_{2}]$ then $V\hat{\mathfrak{h}}_{1}\subseteq g_{2}[y_{2}]\subseteq g_{2}[x_{2}]^{*}=\mathfrak{a}_{2}$ and we are done. So suppose $x_{2}]=[y_{2}]$.
    1. (a)
      Suppose $\mathfrak{h}_{1}\subseteq g_{2}\mathfrak{h}_{0}$. Then $\mathfrak{h}_{1}\subseteq g_{2}[x_{2}]$, so any path from $\mathfrak{a}_{1}\cap\mathfrak{h}_{1}$ to $\mathfrak{a}_{2}=g_{2}[x_{2}]^{*}$ must first cross $\hat{\mathfrak{h}}_{1}$. But $\mathfrak{a}_{1}\cap\mathfrak{h}_{1}\not\preceq\hat{\mathfrak{h}}_{1}$ (indeed this holds for any $(\mathfrak{a},\mathfrak{h})\in\mathcal{P}$), so $\mathfrak{a}_{1}\not\preceq\mathfrak{a}_{2}$, contradicting $(\mathfrak{a}_{1},\mathfrak{h}_{1})\leq(\mathfrak{a}_{2},\mathfrak{h}_{2})$.
    2. (b)
      Suppose $\mathfrak{h}_{1}^{*}\subseteq g_{2}\mathfrak{h}_{0}$. If $V\mathfrak{h}_{1}-V\hat{\mathfrak{h}}_{1}\subseteq\mathfrak{a}_{1}$ then $\mathfrak{a}_{2}=g_{2}[x_{2}]^{*}\subseteq\mathfrak{a}_{1}$, and, combined with $\mathfrak{h}_{1}\not\subseteq\mathfrak{h}_{2}$, this contradicts $(\mathfrak{a}_{1},\mathfrak{h}_{1})\leq(\mathfrak{a}_{2},\mathfrak{h}_{2})$. So $V\mathfrak{h}_{1}-V\hat{\mathfrak{h}}_{1}\not\subseteq\mathfrak{a}_{1}$, and $(\mathfrak{a}_{1},\mathfrak{h}_{1})$ takes the form $(g_{1}[x_{1}],g_{1}\mathfrak{h}_{0})$. By Lemma 4.14 there is $y_{1}]\in\mathcal{M}_{0}$ with $V\mathfrak{h}_{2}=g_{2}V\mathfrak{h}_{0}^{*}\subseteq g_{1}[y_{1}]$. By Lemma 4.15, we have $\mathfrak{a}_{2}=g_{2}[x_{2}]^{*}\preceq g_{1}[y_{1}]$. If $x_{1}]=[y_{1}]$ then $\mathfrak{a}_{2}\preceq\mathfrak{a}_{1}$, and we again contradict $(\mathfrak{a}_{1},\mathfrak{h}_{1})\leq(\mathfrak{a}_{2},\mathfrak{h}_{2})$. So $x_{1}]\neq[y_{1}]$. But then $g_{1}[x_{1}]=\mathfrak{a}_{1}\preceq\mathfrak{a}_{2}\preceq g_{1}[y_{1}]\subseteq g_{1}[x_{1}]^{*}$ and we contradict Lemma 4.16.∎
Lemma 4.23.
There is a constant $K>0$ such that, for any $(\mathfrak{a},\mathfrak{h})\in\mathcal{P}$, any edge $e$ going from $\mathfrak{a}$ to $VX-\mathfrak{a}$ is contained in the $K$-neighborhood of $\hat{\mathfrak{h}}$.
Proof.
Let $K\geq 1$ be such that $\delta$ is contained in the $K$-neighborhood of $\hat{\mathfrak{h}}_{0}$. If $(\mathfrak{a},\mathfrak{h})$ takes the form $(V\mathfrak{h},\mathfrak{h})$ with $\mathfrak{h}\in\mathcal{H}^{\#}$ then $e$ must go from $\hat{\mathfrak{h}}$ to $X-\mathfrak{h}$, so $e$ is certainly in the $K$-neighborhood of $\hat{\mathfrak{h}}$. If $(\mathfrak{a},\mathfrak{h})$ takes the form $(g[x],g\mathfrak{h}_{0})$, then $e$ either goes from $g\hat{\mathfrak{h}}_{0}$ to $X-g\mathfrak{h}_{0}$ or it goes from $g[x]$ to $g[y]$ for some $x\nsim y\in V\mathfrak{h}_{0}$. In the former case, $e$ is certainly in the $K$-neighborhood of $\hat{\mathfrak{h}}=g\hat{\mathfrak{h}}_{0}$. In the latter case, $e$ must belong to a set $gg_{0}\delta$ with $g_{0}\in G_{0}$, so $e$ is in the $K$-neighborhood of $\hat{\mathfrak{h}}=g\hat{\mathfrak{h}}_{0}$ by definition of $K$. If $(\mathfrak{a},\mathfrak{h})$ takes the form $(g[x]^{*},g\mathfrak{h}_{0}^{*})$ then $\mathfrak{a}$ is the complement in $VX$ of $g[x]$, so the argument works the same as the previous case. ∎
Lemma 4.24.
Let $(\mathfrak{a}_{1},\mathfrak{h}_{1})\leq(\mathfrak{a}_{2},\mathfrak{h}_{2})$ in $\mathcal{P}$. For each $\mathfrak{h}_{3}\in\mathcal{H}(X)$ there is at most one $(\mathfrak{a}_{3},\mathfrak{h}_{3})\in\mathcal{P}$ with $(\mathfrak{a}_{1},\mathfrak{h}_{1})\leq(\mathfrak{a}_{3},\mathfrak{h}_{3})\leq(\mathfrak{a}_{2},\mathfrak{h}_{2})$.
Proof.
If $\mathfrak{h}_{3}\in\mathcal{H}^{\#}$ then we must have $\mathfrak{a}_{3}=V\mathfrak{h}_{3}$. Now suppose $\mathfrak{h}_{3}=g\mathfrak{h}_{0}$. Suppose there are two choices for $\mathfrak{a}_{3}$ that work. This means there are distinct $x],[y]\in\mathcal{M}_{0}$ with $(g[x],g\mathfrak{h}_{0}),(g[y],g\mathfrak{h}_{0})$ both greater than $(\mathfrak{a}_{1},\mathfrak{h}_{1})$ and both less than $(\mathfrak{a}_{2},\mathfrak{h}_{2})$. By Lemma 4.18, we have $(g[x],g\mathfrak{h}_{0})\leq(g[y]^{*},g\mathfrak{h}_{0}^{*})=(g[y],g\mathfrak{h}_{0})^{*}$. Also $(\mathfrak{a}_{1},\mathfrak{h}_{1})\leq(g[y],g\mathfrak{h}_{0})$ implies $(g[y],g\mathfrak{h}_{0})^{*}\leq(\mathfrak{a}_{1},\mathfrak{h}_{1})^{*}$. Putting this together we get $(\mathfrak{a}_{1},\mathfrak{h}_{1})\leq(g[x],g\mathfrak{h}_{0})\leq(\mathfrak{a}_{1},\mathfrak{h}_{1})^{*}$, which contradicts $\mathcal{P}$ being a pocset (Definition 3(1)). So there is at most one choice for $\mathfrak{a}_{3}$ that works. Finally, the case $\mathfrak{h}_{3}=g\mathfrak{h}_{0}^{*}$ can be reduced to the case $\mathfrak{h}_{3}=g\mathfrak{h}_{0}$ by applying the involution * of $\mathcal{P}$. ∎
Lemma 4.25.
$\mathcal{P}$ is discrete.
Proof.
[…]
@@ 4.7 The tree $T^{\prime}$
Let $T^{\prime}$ be the cubing of $\mathcal{P}$. We know that $\mathcal{P}$ is discrete by Lemma 4.25 and it has no pair of transverse elements by Lemma 4.21, hence Proposition 3.2 tells us that the cubing $T^{\prime}$ is well-defined and is a tree. Moreover, the action of $G$ on $\mathcal{P}$ induces an action of $G$ on $T^{\prime}$. The edges in $T^{\prime}$ correspond to pairs $\{(\mathfrak{a},\mathfrak{h}),(\mathfrak{a},\mathfrak{h})^{*}\}\subseteq\mathcal{P}$, and, since $(\mathfrak{a},\mathfrak{h})^{*}$ always takes the form $(\mathfrak{a}^{\prime},\mathfrak{h}^{*})$, an edge inversion in $T^{\prime}$ would mean there is an element $g\in G$ with $g\mathfrak{h}=\mathfrak{h}^{*}$. But that would give rise to an edge inversion in $T$, which doesn’t happen, hence $G$ acts on $T^{\prime}$ without edge inversions, i.e. $G\curvearrowright T^{\prime}$ is a splitting. We may assume that the action $G\curvearrowright T^{\prime}$ is minimal since we can always replace $T^{\prime}$ with a minimal $G$-invariant subtree. The description of edges in $T^{\prime}$ also makes it clear that every edge stabilizer of $T^{\prime}$ is contained in an edge stabilizer of $T$.
@@
$\phi(v):=$ $\{\mathfrak{h}\in\mathcal{H}^{\#}\mid(V\mathfrak{h},\mathfrak{h})\in v\}$
@@
Let’s first prove completeness of $\phi(v)$. For $\mathfrak{h}\in\mathcal{H}^{\#}$, exactly one of $(V\mathfrak{h},\mathfrak{h})$ and $(V\mathfrak{h}^{*},\mathfrak{h}^{*})$ is in $v$ by completeness of $v$, so exactly one of $\mathfrak{h}$ and $\mathfrak{h}^{*}$ is in $\phi(v)$. For $g\in G$ and $x]\in\mathcal{M}_{0}$, exactly one of $(g[x],g\mathfrak{h}_{0})$ and $(g[x]^{*},g\mathfrak{h}_{0}^{*})$ is in $v$ by completeness of $v$, so for each $g\in G$ we have exactly one of $g\mathfrak{h}_{0}$ and $g\mathfrak{h}_{0}^{*}$ in $\phi(v)$.
Before showing that $\phi(v)$ is consistent and DCC, we make the following observation, which follows from Lemmas 4.14 and 4.15: if $\mathfrak{h}_{1}\subsetneq\mathfrak{h}_{2}$ in $\mathcal{H}(X)$ then
@@
    if $\mathfrak{h}_{1}\in\mathcal{H}^{\#}$ then $\mathfrak{a}_{1}=V\mathfrak{h}_{1}$,
@@
    if $\mathfrak{h}_{1}=g_{1}\mathfrak{h}_{0}$ then $\mathfrak{a}_{1}=g_{1}[x]$ and (16) holds for all choices of $x]\in\mathcal{M}_{0}$,
@@
    if $\mathfrak{h}_{1}=g_{1}\mathfrak{h}_{0}^{*}$ then $\mathfrak{a}_{1}=g_{1}[x]^{*}$ for the unique $x]\in\mathcal{M}_{0}$ with $V\mathfrak{h}_{2}^{*}\subseteq g_{1}[x]$,
@@
    if $\mathfrak{h}_{2}\in\mathcal{H}^{\#}$ then $\mathfrak{a}_{2}=V\mathfrak{h}_{2}$,
@@
    if $\mathfrak{h}_{2}=g_{2}\mathfrak{h}_{0}$ then $\mathfrak{a}_{2}=g_{2}[y]$ for the unique $y]\in\mathcal{M}_{0}$ with $V\mathfrak{h}_{1}\subseteq g_{2}[y]$,
@@
    if $\mathfrak{h}_{2}=g_{2}\mathfrak{h}_{0}^{*}$ then $\mathfrak{a}_{2}=g_{2}[y]^{*}$ and (16) holds for all choices of $y]\in\mathcal{M}_{0}$.
We now prove consistency of $\phi(v)$. If $\mathfrak{h}_{1}\subsetneq\mathfrak{h}_{2}$ and $\mathfrak{h}_{1}\in\phi(v)$ then $(\mathfrak{a}_{1},\mathfrak{h}_{1})\in v$, where $(\mathfrak{a}_{1},\mathfrak{h}_{1})$ is given by the appropriate case from (1)(3) (and in case (2) we only have $(g_{1}[x],g_{1}\mathfrak{h}_{0})\in v$ for one $x]\in\mathcal{M}_{0}$). Consistency of $v$ then implies that $(\mathfrak{a}_{2},\mathfrak{h}_{2})\in v$, where $(\mathfrak{a}_{2},\mathfrak{h}_{2})$ is given by the appropriate case from (4)(6) (and in case (6) we have $(g_{2}[y]^{*},g_{2}\mathfrak{h}_{0}^{*})\in v$ for all $y]\in\mathcal{M}_{0}$). It follows that $\mathfrak{h}_{2}\in\phi(v)$.
Finally, we prove that $\phi(v)$ is DCC. Suppose for contradiction that we have an infinite descending chain $\mathfrak{h}_{1}\supsetneq\mathfrak{h}_{2}\supsetneq\mathfrak{h}_{3}\supsetneq\dots$ of distinct elements in $\phi(v)$. Then we can turn this into an infinite descending chain
@@
    if $\mathfrak{h}_{i}\in\mathcal{H}^{\#}$ then $\mathfrak{a}_{i}=V\mathfrak{h}_{i}$,
@@
    if $\mathfrak{h}_{i}=g\mathfrak{h}_{0}$ then $\mathfrak{a}_{i}=g[x]$ for the unique $x]\in\mathcal{M}_{0}$ with $V\mathfrak{h}_{i+1}\subseteq g[x]$,
@@
    if $\mathfrak{h}_{i}=g\mathfrak{h}_{0}^{*}$ then $\mathfrak{a}_{i}=g[x]^{*}$ for the unique $x]\in\mathcal{M}_{0}$ with $V\mathfrak{h}_{i-1}^{*}\subseteq g[x]$ (or any $x]\in\mathcal{M}_{0}$ if $i=1$).
@@
Recall that we have a wallspace $(V\mathfrak{h}_{0},\mathcal{P}_{0})$ with cubing $T_{0}$, and an associated map $\lambda:V\mathfrak{h}_{0}\to VT_{0}$. Moreover, it follows from the constructions that $\lambda$ is $G_{0}$-equivariant. We observed that the $\sim$-classes $[x]$ are precisely the non-empty fibers of $\lambda$, so in particular we get a well-defined injective $G_{0}$-equivariant map $\mathcal{M}_{0}\to VT_{0}$ given by $x]\mapsto\lambda(x)$.
@@
  • For $\mathfrak{h}\in\mathcal{H}^{\#}$, we have an edge $e^{\prime}=\{(V\mathfrak{h},\mathfrak{h}),(V\mathfrak{h}^{*},\mathfrak{h}^{*})\}\in ET^{\prime}$ and also an edge $e\in ET$. In this case $e\notin G\cdot e_{0}$, so the tree $T_{e}$ is just a single vertex, and we define $\tau(e^{\prime})$ to be this vertex. The maps $\mathcal{H}^{\#}\to ET$ and $\mathcal{H}^{\#}\to ET^{\prime}$ are $G$-equivariant, so $\tau(ge^{\prime})=g\cdot\tau(e^{\prime})$ for all $g\in G$.
@@
    with $i\in\Omega$ and $x]\in\mathcal{M}_{0}$. We then define
@@
Finally, the map $\tau:ET^{\prime}\to\sqcup_{e}VT_{e}$ is injective because the map $\mathcal{M}_{0}\to VT_{0}$ given by $x]\mapsto\lambda(x)$ is injective. ∎
@@
Recall that $e_{0}\in ET$ is the edge corresponding to the halfspace $\mathfrak{h}_{0}$, in particular $G_{0}=G_{e_{0}}$. The vertex $v_{0}\in VT$ will be the endpoint of $e_{0}$ that is contained in the halfspace $p(\mathfrak{h}_{0})\in\mathcal{H}(T)$. The vertex stabilizer $G_{v_{0}}$ fixes no edges in $T$, so there exists $g\in G_{v_{0}}-G_{e_{0}}$. Then $g\mathfrak{h}_{0}^{*}\subsetneq\mathfrak{h}_{0}$. By Lemma 4.15, there exist $x],[y]\in\mathcal{M}_{0}$ with $g[y]^{*}\preceq[x]$, so $(g[y]^{*},g\mathfrak{h}_{0}^{*})\leq([x],\mathfrak{h}_{0})$ in $\mathcal{P}$. Recall that we have a $G_{0}$-equivariant map $\mathcal{M}_{0}\to T_{0}$ (as in the proof of Lemma 4.27), and $G_{0}$ has no fixed point in $T_{0}$ by Lemma 4.13, so the $G_{0}$-orbits in $\mathcal{M}_{0}$ are all infinite. In particular, there exists $g_{0}\in G_{0}$ with $g_{0}[x]\neq[y]$. Then $(g_{0}[x],\mathfrak{h}_{0})\leq([y]^{*},\mathfrak{h}_{0}^{*})$ by Lemma 4.18. Putting this all together, we get